# ALCM: Autonomous LLM-Augmented Causal Discovery Framework

### v1

@@ 1 Introduction
Within the literature, numerous studies have contributed significantly to the development of a variety of efficient causal discovery algorithms aimed at uncovering the underlying causal structure from observational data. This body of research can be broadly categorized into two main groups: conventional data-driven causal discovery algorithms and those based on LLMs. Conventional causal discovery algorithms focus on learning the causal graph from samples of the joint probability distribution of observational data. They utilize various statistical techniques, including conditional independence tests, machine learning approaches, generative models, deep learning methodologies, and reinforcement learning strategies [35] to understand the joint distribution of observed variables and extract the causal connections among them. Subsequently, these algorithms assess how well the candidate causal graph aligns with the data [48, 14, 13].
@@
In this paper, we present an LLM-powered causal discovery framework–ALCM: a multi-component Autonomous LLM-Augmented Causal Discovery Framework. ALCM proposes a synergized reasoning method and entails three components: causal structure learning, causal wrapper, and LLM-driven refiner components to generate more accurate and robust causal graphs. ALCM is engineered to autonomously untangle causal structures by deciphering those causal relations embedded in observational data. ALCM capitalizes on observed data, data-driven causal reasoning algorithms, and the implicit knowledge embedded in LLMs to optimize and streamline the entire causal reasoning process. This approach aims to establish a more robust, applicable, and reliable foundation for causal reasoning and estimation as well. We conduct a comprehensive performance evaluation of ALCM, employing LLMs and assessing their capabilities on widely recognized benchmarks [36, 42]. We compare our framework with conventional causal discovery algorithms and LLMs prompting. Furthermore, we implement an automatic pipeline for making the causal discovery an automatic task.
@@ 3 Proposed Framework
In this section, we present ALCM, an advanced causal discovery framework aimed to leverage the combined strengths of traditional causal discovery algorithms and LLMs. ALCM provide an automated pipeline constructing a comprehensive causal graph, refining it, and incorporating previously overlooked insights to enrich the resulting causal model. This integration aims to utilize the precision of conventional causal discovery algorithms in identifying data relationships, while also enhancing and validating these findings with insights from LLMs. Fig. 1 indicates an overview of the ALCM framework. The algorithmic perspective of the ALCM framework is detailed in Algorithm 1. The ALCM framework includes three principal components: Causal Structure Learning, Causal Wrapper, and LLM-driven Refiner. To clarify the functionality and definitions of the framework, we present and exemplify these components in the following.
@@
1:Observed dataset, $O$; Contextual Causal Information, $C$; Metadata, $M$
2:Causal DAG, $DAG$
3:Initialize and run the selected data-driven causal discovery algorithms CD, $G_{\text{i}}\leftarrow CD(O)$
4:Generate the causal prompt by injecting C and M
5:for each $z=(e_{i},e_{j})$ in $G_{\text{i}}$ do
6:     if $z$ is validated by LLM-Driven Refiner then
7:         $G_{\text{i}}\leftarrow G_{\text{i}}\cup\emptyset$
8:     end if
9:     if $z$ orientation is revised by LLM-Driven Refiner then
10:         $G_{\text{i}}\leftarrow z^{\prime}\cup G_{\text{i}}$
11:     end if
12:     if $z$ is removed by LLM-Driven Refiner then
13:         $G_{\text{i}}\leftarrow G_{\text{i}}-z^{\prime}$
14:     end if
15:     if a new $z^{\prime\prime}$ is added by LLM-Driven Refiner then
16:         $G_{\text{i}}\leftarrow z^{\prime\prime}\cup G_{\text{i}}$
17:     end if
18:end for
19:return $G_{\text{i}}$
@@ 3.1 Causal Structure Learning
The Causal Structure Learning component is our data-driven conventional causal discovery component. It receives a dataset as its input and generates the initial causal graphs from the dataset. This component uncovers causal insights by analyzing purely observational data, and it develops graphical structures that can be interpreted causally. This component directly influences the accuracy and reliability of both the final causal graph and future causal inferences drawn from the data.
Conventional causal discovery algorithms can be leveraged in this component. For instance, we can implement conventional causal discovery algorithms, including Peter-Clark (PC) [38] and Linear Non-Gaussian Acyclic Model (LiNGAM) [37] to discern the probabilistic dependencies and independencies among variables. These algorithms are selected based on their proven efficacy in uncovering intricate causal relationships within complex data.
For the implementation part, we implement PC conventional causal discovery algorithms; moreover, we propose a hybrid method combining the PC and LiNGAM algorithms to identify causal relationships effectively. Building on this foundation, the causal structure learning component crafts an initial causal graph, encapsulating the potential causal linkages derived from the datasets. Finally, this generated causal graph is then relayed to the Causal Wrapper component for further processing.
@@ 3.2 Causal Wrapper
The Causal Wrapper component serves as a critical intermediary or bridge between the causal structure learning and LLM-driven refiner components. This component encapsulates and translates the raw, initial causal graph into a series of contextual, causal-aware prompts (i.e., causal prompts). These prompts are fed to the LLM-driven refiner. The primary aim of these causal prompts is to act as guides for the LLM-driven refiner, aiding it in comprehending the initial causal graph. Furthermore, these causal prompts direct the LLM-driven refiner to identify and integrate the relevant and updated causal background knowledge to make the solution more suited to the specific causal discovery problem at hand. Given these reasons, this prompting strategy ensures that the final causal graph is not only precise but also robust and reflective of the underlying causal mechanisms within the dataset.
Equation 1 shows our causal-aware prompting strategy by infusing the context of problem and metadata information into the prompts. This prompting strategy was inspired from an effort by Kim et al. [21]. They demonstrated that contextual information is important in boosting the overall performance of LLMs’ responses.
@@
Figure 2 illustrates an example of the causal wrapper’s functionality. Additionally, the output can incorporate supplementary reasoning and confidence levels for the generated response. For instance, a simple instruction can prompt the LLM-driven refiner to engage in step-by-step reasoning, employing a Chain-of-Thought (CoT) approach [45]. Moreover, it can request the LLM to indicate its level of confidence or likelihood regarding the generated output, using either a log-likelihood value or a confidence level. Once these causal prompts are generated, they are dispatched to the LLM-driven refiner component. This method ensures that the ALCM framework optimally utilizes LLMs for uncovering, refining, and validating causal relationships, thereby advancing the field of causal discovery with a high level of accuracy.
@@ 3.3 LLM-driven Refiner
The LLM-driven Refiner leverages advanced language models in the refinement and evaluation of causal graphs. This component receives a series of intricately designed, contextual causal prompts from the causal wrapper component, which serve as a nuanced guide for its operations. Its core function is to assess, refine, and potentially augment the initial causal graph by evaluating the causal edges/nodes, and, where necessary, adding or removing nodes and edges to represent the underlying causal mechanisms better.
@@ 4.1 Implementation 1 (ALCM-PC)
For the first implementation, we select PC causal discovery algorithm. The PC algorithm is renowned for its robustness in dealing with large datasets and its ability to infer causal structures through conditional independence tests, making it highly efficient in uncovering complex causal networks. For the causal wrapper component, we utilize causal prompt. We illustrate one example of our prompt in Figure 3. For LLM-driven refiner, we exploit OpenAI GPT-4 [29, 3] in our pipeline.
@@ 4.2 Implementation 2 (ALCM-Hybrid)
For the second implementation, we leverage a hybrid approach (which outlined in Section 3.1) including PC [38] and LiNGAM [37] algorithms due to their complementary strengths and proven effectiveness in identifying causal relationships. This hybrid method utilizes a majority vote mechanism for identifying common causal edges recognized by both algorithms. For edges that are uniquely identified by only one algorithm and not common to both, we introduce an extra step by employing LLMs as a decisive judge. This entails presenting these edges to LLMs to ascertain their potential as causal links based on contextual understanding and causal reasoning capabilities. If affirmed, these edges are added to an augmented set of causal connections. The causal wrapper component applies the causal prompt template and the result is sent to the LLM-driven refiner. The LLM-driven refiner (OpenAI model) Component evaluates, refines, and enhances the causal graph to produce a final, enhanced causal structure.

### v2

@@ 1 Introduction
Within the literature, numerous studies have contributed significantly to the development of a variety of efficient causal discovery algorithms aimed at uncovering the underlying causal structure from observational data. This body of research can be broadly categorized into two main groups: conventional data-driven causal discovery algorithms and those based on LLMs [44]. Conventional causal discovery algorithms focus on learning the causal graph from samples of the joint probability distribution of observational data. They utilize various statistical techniques, including conditional independence tests, machine learning approaches, generative models, deep learning methodologies, and reinforcement learning strategies [35] to understand the joint distribution of observed variables and extract the causal connections among them. Subsequently, these algorithms assess how well the candidate causal graph aligns with the data [49, 14, 13].
@@
In this paper, we present an LLM-powered causal discovery framework–ALCM: a multi-component Autonomous LLM-Augmented Causal Discovery Framework. ALCM proposes a synergized reasoning method and entails three components: causal structure learning, causal wrapper, and LLM-driven refiner components to generate more accurate and robust causal graphs. ALCM is engineered to autonomously untangle causal structures by deciphering those causal relations embedded in observational data. ALCM capitalizes on observed data, data-driven causal reasoning algorithms, and the implicit knowledge embedded in LLMs to optimize and streamline the entire causal reasoning process. This approach aims to establish a more robust, applicable, and reliable foundation for causal reasoning and estimation as well. We conduct a comprehensive performance evaluation of ALCM, employing LLMs and assessing their capabilities on widely recognized benchmarks [36, 42]. We compare our framework with conventional causal discovery algorithms and LLMs prompting. Furthermore, we implement an automatic pipeline for making the causal discovery an automatic task. Our contributions are as follows:
Our contributions in this work are as follows:
  • Unified Framework for Enhanced Causal Discovery: We introduce the ALCM framework that synergistically integrates the strengths of conventional data-driven causal discovery (CCD) methods and Large Language Models (LLMs) to overcome the limitations of individual approaches by generating accurate, interpretable, and comprehensive causal graphs.
  • Dynamic, Scalable, and Autonomous Operations: ALCM demonstrates adaptability to dynamic data environments, autonomous operation without domain expertise dependency, and scalability across unseen datasets.
  • Improved Predictive Precision and Reliability: Leveraging the contextual reasoning capabilities of LLMs, the framework refines causal relationships with algorithmic rigor.
  • Comprehensive Graph Representation and Explainability: ALCM provides a fully automated pipeline for constructing and refining causal graphs. It ensures interpretability and explainability of results.
  • Benchmarking and Performance Validation: We extensively evaluate ALCM across multiple benchmark datasets, demonstrating its superior performance compared to the related works.
This work advances the field of causal discovery by demonstrating the transformative potential of a unified framework that synthesizes the complementary capabilities of CCD and LLM-based approaches.
@@ 3 Proposed Framework
In this section, we present ALCM, an advanced causal discovery framework aimed to leverage the combined strengths of traditional causal discovery algorithms and LLMs. ALCM provide an automated pipeline constructing a comprehensive causal graph, refining it, and incorporating previously overlooked insights to enrich the resulting causal model. This integration aims to utilize the precision of conventional causal discovery algorithms in identifying data relationships, while also enhancing and validating these findings with insights from LLMs. Fig. 1 indicates an overview of the ALCM framework. The algorithmic perspective of the ALCM framework is detailed in Algorithm 1. The ALCM framework includes three principal components: Causal Structure Learning, Causal Wrapper, and LLM-driven Refiner. These components interact iteratively to enhance the causal discovery process, to clarify the functionality and definitions of the framework, we present and exemplify these components and their interconnections in the following.
@@
Require: Observed dataset, $O$; Contextual Causal Information, $C$; Metadata, $M$
Ensure: Causal DAG, $DAG$
1:Initialize and run the selected data-driven causal discovery algorithm, $CD$, to generate an initial causal graph $G_{i}\leftarrow CD(O)$triangleright Step 1: Generate the initial causal graph from observational data.
2:Generate the causal prompt by injecting $C$ and $M$triangleright Step 2: Prepare contextual information for refinement by LLM.
3:for each edge $z=(e_{i},e_{j})$ in $G_{i}$ do
4:     if $z$ is validated by LLM-Driven Refiner then $\triangleright$ Step 3: Validate the existence of the causal edge using the LLM.
5:         $G_{i}\leftarrow G_{i}\cup\emptyset$triangleright If valid, retain the edge.
6:     end if
7:     if $z$ orientation is revised by LLM-Driven Refiner then $\triangleright$ Step 4: Adjust the edge direction if required.
8:         $G_{i}\leftarrow z^{\prime}\cup G_{i}$
9:     end if
10:     if $z$ is removed by LLM-Driven Refiner then $\triangleright$ Step 5: Remove the edge if deemed invalid.
11:         $G_{i}\leftarrow G_{i}-z^{\prime}$
12:     end if
13:     if a new edge $z^{\prime\prime}$ is added by LLM-Driven Refiner then $\triangleright$ Step 6: Add a new edge identified by the LLM.
14:         $G_{i}\leftarrow z^{\prime\prime}\cup G_{i}$
15:     end if
16:end for
17:Return $G_{i}$
@@ 3.1 Causal Structure Learning
The Causal Structure Learning component serves as the foundational data-driven module of the ALCM framework, responsible for generating the initial causal graph from observational datasets. This component identifies causal relationships among variables by analyzing probabilistic dependencies and independencies in the data, leveraging well-established conventional causal discovery methods. These methods typically infer causal graphs by estimating relationships between variables (nodes) and their potential causal links (edges), using statistical tests to determine conditional independencies and distinguish direct relationships from those mediated by other variables. Additionally, orientation rules are applied to establish causal directions, guided by assumptions such as acyclicity or specific properties of the data distribution. Optimization techniques, including scoring functions-based methods (e.g., Bayesian Information Criterion (BIC) or likelihood-based scores), might be used to refine graph structures and select the most plausible causal model. To enhance robustness, the causal structure learning component integrates complementary methods to account for diverse data characteristics, such as linearity, non-linearity, or Gaussianity. The output of this component is an initial causal graph encapsulating key variables (nodes) and their inferred causal relationships (edges). The causal structure learning component directly influences the accuracy and reliability of both the final causal graph and future causal inferences drawn from the data.
For the implementation, we utilize three established causal discovery algorithms—Peter-Clark (PC) [38], Linear Non-Gaussian Acyclic Model (LiNGAM) [37], and Non-combinatorial Optimization via Trace Exponential and Augmented lagRangian for Structure learning (NOTEARS) [51]—each chosen for their distinct strengths and complementary characteristics.
The PC algorithm leverages conditional independence (CI) tests to construct a graph’s skeleton and identify v-structures, offering a robust framework for causal inference when the underlying relationships can be uncovered through probabilistic dependencies. Its ability to handle discrete and continuous variables effectively makes it a reliable choice for datasets where independence testing plays a central role. LiNGAM, on the other hand, excels at uncovering linear causal relationships in datasets with non-Gaussian distributions. By utilizing Independent Component Analysis (ICA) for causal ordering, LiNGAM demonstrates superior performance in disentangling complex linear interactions. Its focus on exploiting the statistical properties of non-Gaussianity ensures that causal directions are accurately inferred, even in the presence of latent confounders. Complementing these approaches, NOTEARS offers a novel optimization-based framework that reformulates the combinatorial problem of DAG discovery into a continuous optimization task. By incorporating an acyclicity constraint into its objective function, NOTEARS efficiently learns causal structures while maintaining scalability to larger datasets. Its gradient-based methodology makes it particularly adept at handling high-dimensional data with intricate causal dependencies.
Building on the unique strengths of these three algorithms, we propose a hybrid method that combines PC, LiNGAM, and NOTEARS within a unified framework. This hybrid approach employs a majority-weighted voting mechanism to leverage the individual advantages of each algorithm dynamically. The weighting is determined based on their relative performance on specific datasets, ensuring that the final causal graph benefits from their collective expertise. This integration enhances the robustness and reliability of the causal discovery process, allowing the hybrid method to adapt to diverse data characteristics.
The causal structure learning component synthesizes an initial causal graph by combining the outputs of these algorithms, encapsulating the potential causal linkages identified from the dataset. This graph, which represents the key variables (nodes) and their inferred causal relationships (edges), is subsequently passed to the Causal Wrapper component for further contextualization and refinement, enabling downstream tasks to operate on a well-defined and accurate causal structure.
@@ 3.2 Causal Wrapper
The Causal Wrapper component serves as a critical intermediary or bridge between the causal structure learning and LLM-driven refiner components. This component encapsulates and translates the raw, initial causal graph into a series of contextual, causal-aware prompts (i.e., causal prompts). These prompts are fed to the LLM-driven refiner. The primary aim of these causal prompts is to act as guides for the LLM-driven refiner, aiding it in comprehending the initial causal graph. Furthermore, these causal prompts direct the LLM-driven refiner to identify and integrate the relevant and updated causal background knowledge to make the solution more suited to the specific causal discovery problem at hand. Given these reasons, this prompting strategy ensures that the final causal graph is not only precise, but also robust and reflective of the underlying causal mechanisms within the dataset.
Equation 1 shows our causal-aware prompting strategy by infusing the context of problem and metadata information into the prompts. This prompting strategy was inspired by an effort by Kim et al. [21]. They demonstrated that contextual information is important in boosting the overall performance of LLMs’ responses.
@@
Figure 2 illustrates an example of the causal wrapper’s functionality. The causal structure learning component generates the initial causal graph by applying conventional causal discovery algorithms, such as Peter-Clark (PC), LiNGAM, or NOTEARS. These algorithms analyze the input observational dataset to uncover key variables (nodes) and their probabilistic dependencies, forming the skeleton of the initial graph. The nodes in this graph represent significant variables derived from the dataset, while the directed edges illustrate potential causal relationships. This foundational graph is then passed to the causal wrapper and subsequently refined by the LLM-driven refiner. Notably, the output can incorporate supplementary reasoning and confidence metrics to enhance interpretability. For instance, a simple instruction can prompt the LLM-driven refiner to engage in step-by-step reasoning using a Chain-of-Thought (CoT) approach [46]. Additionally, it can request the LLM to quantify its confidence level or provide a likelihood estimate for the generated outputs, using log-likelihood values or confidence percentages.
These prompts are critical for ensuring that the LLM comprehends the initial graph and integrates relevant causal knowledge effectively. For example, they direct the LLM to identify hidden causal relationships or validate existing edges by leveraging its contextual reasoning capabilities. This interaction facilitates the enhancement of the initial graph into a more accurate and robust representation of the underlying causal mechanisms. Once these causal prompts are generated, they are dispatched to the LLM-driven refiner component. This method ensures that the ALCM framework optimally utilizes LLMs for uncovering, refining, and validating causal relationships, thereby advancing the field of causal discovery with a high level of accuracy.
@@ 3.3 LLM-driven Refiner
The LLM-driven Refiner leverages advanced language models in the refinement and evaluation of causal graphs. This component receives a series of intricately designed, contextual causal prompts from the causal wrapper component, which serve as a nuanced guide for its operations.
The LLM-driven Refiner evaluates each edge and node in the graph by applying advanced reasoning capabilities of LLMs (e.g., GPT-4). The process involves:
  1. 1.
    Assessing the validity of existing causal relationships using contextual knowledge.
  2. 2.
    Detecting and integrating hidden causal relationships by reasoning over unobserved variables.
  3. 3.
    Reorienting or removing edges that do not align with domain knowledge or probabilistic dependencies.
  4. 4.
    Assigning confidence scores or likelihood estimates to refined relationships, ensuring interpretability and reliability.
The LLM-driven Refiner verifies hidden causal relationships by leveraging advanced capabilities of LLMs to assess, validate, and refine the initial causal graph. This process begins with the causal prompts generated by the Causal Wrapper, which provide the LLM with explicit instructions, contextual metadata, and domain-specific information about the causal relationships. The LLM evaluates each edge in the graph based on its internal knowledge base and the provided context, determining whether the relationship is valid, needs reorientation, or should be removed. Furthermore, the LLM identifies potential hidden relationships by reasoning over unobserved variables and interactions that conventional algorithms may overlook. To ensure accuracy, the refined relationships are accompanied by confidence scores or likelihood estimates, enabling a structured and interpretable refinement process. Finally, these refined graphs are validated through cross-referencing with up-to-date domain knowledge, ensuring that the final output is both accurate and comprehensive.
@@
3.4 Interactions Between Components
The Causal Structure Learning component generates the initial graph, providing a data-driven foundation. The Causal Wrapper transforms this graph into contextualized prompts, enabling the LLM to reason about relationships in a guided and structured manner. The LLM-driven Refiner refines and validates the graph, identifying hidden relationships and ensuring alignment with external knowledge. This iterative feedback loop ensures that the final causal graph is both accurate and interpretable, addressing limitations in traditional causal discovery methods and leveraging the strengths of LLMs.
@@ 4.1 Implementation 1 (ALCM-PC)
For the first implementation, we select the PC causal discovery algorithm for its robustness in handling large datasets and its efficiency in inferring causal structures through conditional independence (CI) tests. The algorithm constructs an undirected graph and iteratively removes edges by testing CI between variable pairs, conditioned on subsets of other variables, a process known as skeleton discovery. It then orients edges by detecting v-structures (triplets of nodes with specific dependency patterns) and ensuring acyclicity, resulting in a DAG that represents the causal relationships. The PC algorithm’s ability to prune unnecessary connections makes it particularly effective for high-dimensional datasets, balancing computational efficiency with accuracy. For the causal wrapper component, we utilize causal prompt. We illustrate one example of our prompt in Figure 3. For LLM-driven refiner, we exploit OpenAI GPT-4 [29, 3] in our pipeline.
@@ 4.2 Implementation 2 (ALCM-Hybrid)
For the second implementation, we present a hybrid causal discovery approach that integrates the strengths of three leading methods: Peter-Clark (PC) [38], LiNGAM [37], and NOTEARS [51]. These methods address different aspects of causal structure learning, and their combination provides a robust and accurate framework for causal discovery.
The PC method employs conditional independence (CI) tests to iteratively construct a causal graph by building its skeleton and identifying v-structures. This method is particularly effective for datasets with mixed discrete and continuous variables and excels in capturing probabilistic dependencies. Its iterative and constraint-based nature ensures computational efficiency, even in high-dimensional settings. In contrast, LiNGAM is specifically designed to uncover linear causal relationships in datasets with non-Gaussian distributions. By leveraging Independent Component Analysis (ICA), LiNGAM accurately identifies causal ordering and orients edges with high precision, even in the presence of latent confounders and linear dependencies. NOTEARS complements these approaches by reformulating causal discovery into a continuous optimization problem. By incorporating a differentiable acyclicity constraint, NOTEARS transforms the combinatorial problem of DAG discovery into a solvable optimization task, making it highly effective for datasets with intricate causal dependencies and scalable to high-dimensional data.
To leverage the unique strengths of these methods, we propose a hybrid approach that combines their outputs using dynamically assigned weights. These weights are determined based on a composite score for each method, which captures its performance on a given dataset. The composite score is defined as the difference between the Accuracy and NHD, balancing edge-specific performance and structural alignment with the ground truth. Formally, the composite score for a method is given by:
$\text{Composite}_{\text{method}}=\text{Accuracy}_{\text{method}}-\text{NHD}_{\text{method}}$ (2)
This score accounts for both the overall correctness of edge identification (via Accuracy) and the structural similarity of the causal graph (via NHD), ensuring that methods achieving both accurate and well-aligned graphs are given higher importance. The weights are derived by normalizing the composite scores across all methods:
$W_{\text{method}}=\frac{\text{Composite}_{\text{method}}}{\sum_{\text{all methods}}\text{Composite}_{\text{method}}}$ (3)
where $W_{\text{method}}$ represents the weight assigned to a method, ensuring that the sum of all weights equals one.
To further enhance the adaptability of the hybrid approach, we introduce a neural network-based architecture to dynamically learn these weights based on both method performance metrics and dataset-specific features. The neural network is designed to take as input the composite scores of the methods, along with features such as graph density, node degree distribution, and sparsity. Graph density quantifies how connected the graph is and is defined as the ratio of the number of edges to the maximum possible edges. Node degree distribution describes the variability in the number of connections per node, while sparsity measures the proportion of missing edges compared to a fully connected graph.
The architecture of the neural network consists of three layers: 1. An input layer with nine features, including the composite scores of the methods ($\text{Composite}_{\text{PC}},\text{Composite}_{\text{LiNGAM}},\text{Composite}_{\text{NOTEARS}}$) and six dataset-specific features such as graph density, average node degree, and sparsity. 2. Two hidden layers with 64 and 32 neurons, respectively, each using the Rectified Linear Unit (ReLU) activation function to capture non-linear relationships among the features. 3. An output layer with three neurons (one for each method), using a softmax activation to produce normalized weights for the methods. Formally, the neural network outputs the weights as follows:
$\mathbf{W}=\text{Softmax}(\mathbf{H}_{2}\cdot\mathbf{W}_{o}+\mathbf{b}_{o})$ (4)
where $\mathbf{H}_{2}$ represents the outputs from the second hidden layer, $\mathbf{W}_{o}$ and $\mathbf{b}_{o}$ are the weights and biases of the output layer, and Softmax ensures the weights sum to one.
The neural network is trained on a dataset comprising simulated graphs with varying densities, node degrees, and sparsity levels. For each graph, the outputs of PC, LiNGAM, and NOTEARS are evaluated using Accuracy and NHD, and the composite scores are computed. The ground-truth weights for training are derived by normalizing these composite scores as described in Equation (3). The training objective minimizes the mean squared error (MSE) between the predicted weights and the ground-truth weights. Using the dynamically learned weights, the hybrid approach synthesizes a causal graph by aggregating the outputs of PC, LiNGAM, and NOTEARS. For each edge $e$, the final score is computed as:
$\text{Score}_{e}=\sum_{\text{method}}W_{\text{method}}\cdot\mathbb{I}_{\text{method}(e)}$ (5)
Here, $\mathbb{I}_{\text{method}(e)}$ is an indicator function that equals 1 if the edge $e$ is identified by the method and 0 otherwise. Edges with scores exceeding a predefined threshold are retained in the hybrid causal graph. For edges uniquely identified by only one method, one LLM is employed as a decisive layer. The LLM evaluates these edges based on contextual knowledge and causal reasoning to ensure that only plausible causal links are included. The validated edges are then added to the hybrid graph, enhancing its comprehensiveness and accuracy.
The resulting hybrid causal graph is passed to the Causal Wrapper component, where it is further contextualized and refined using domain-specific templates and LLM-driven reasoning. This ensures that the final causal graph is robust, accurate, and adaptable to diverse data characteristics. By combining the strengths of traditional causal discovery methods with the adaptability of neural networks and the reasoning capabilities of LLMs, the hybrid approach achieves superior performance in causal discovery, setting a new benchmark for accuracy and robustness.